Problem 6-02A a, b1-b2, c (Video)

Martinez Distribution markets CDs of the performing artist Fishe. At the beginning of October, Martinez had in beginning inventory 3,100 of Fishe’s CDs with a unit cost of $7. During October, Martinez made the following purchases of Fishe’s CDs.
Oct. 3 3,875 @ $8 Oct. 19 4,650 @ $10
Oct. 9 5,425 @ $9 Oct. 25 6,200 @ $11

During October, 16,895 units were sold. Martinez uses a periodic inventory system.
Determine the cost of goods available for sale.
Cost of goods available for sale $
Calculate weighted average cost per unit. (Round answer to 2 decimal places, e.g. 2.25.)
Weighted average cost per unit $
Determine (1) the ending inventory and (2) the cost of goods sold under each of the assumed cost flow methods (FIFO, LIFO, and average-cost). Prove the accuracy of the cost of goods sold under the FIFO and LIFO methods. (Round answers to 0 decimal places, e.g. 1,250. Use weighted-average unit cost rounded to 2 decimal places for computations.)

FIFO

LIFO

AVERAGE-COST

The ending inventory $ $ $
The cost of goods sold $ $ $
Which cost flow method results in (1) the highest inventory amount for the balance sheet and (2) the highest cost of goods sold for the income statement?
(1)

FIFOLIFOAverage-cost

results in the highest inventory amount, $ .
(2)

FIFOLIFOAverage-cost

results in the highest cost of goods sold, $ .
Answer #1

Calculate following

Unit Unit Cost Total Cost
Beg Inv 3100 7 21700
Purchase 1 3875 8 31000
Purchase 2 5425 9 48825
Purchase 3 4650 10 46500
Purchase 4 6200 11 68200
Total 23250 216225

a) Cost of goods available for sale = 216225

b) Weighted average cost per unit = 216225/23250 = 9.3

c) Calculate following

FIFO LIFO Average cost
The ending inventory 68200+155*10 = 69750 21700+3255*8 = 47740 6355*9.3 = 59101.5 or 59102
The cost of goods sold 216225-69750 = 146475 216225-47740 = 168485 216225-59101.50 = 157123.50 or 157123

Prove accuracy :

FIFO LIFO Average cost
The ending inventory 68200+155*10 = 69750 21700+3255*8 = 47740 6355*9.3 = 59101.5 or 59102
The cost of goods sold 216225-69750 = 146475 216225-47740 = 168485 216225-59101.50 = 157123.50 or 157123
Cost of goods available for sale 216225 216225 216225

Highest inventory = FIFO

Highest cost of goods sold = LIFO